OpenStudy (anonymous):

Write an equation of the line that passes through (8, 3) and (4, 3).

jimthompson5910 (jim_thompson5910):

Since the y coordinates are both equal to 3, the equation is y = 3 This is a horizontal line passing through 3 on the y-axis

OpenStudy (anonymous):

Will you explain to me how to figure this out?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Because I have several of these questions, and I want to understand how to do them

jimthompson5910 (jim_thompson5910):

sure just post them to the left or you can ask here

OpenStudy (anonymous):

Here is one that I dont understand -- Write the equation of the line that passes through (-4, 12) and (7, 23).

jimthompson5910 (jim_thompson5910):

ok one sec

OpenStudy (anonymous):

k

jimthompson5910 (jim_thompson5910):

step 1 is to find the slope, how do we do that?

OpenStudy (anonymous):

Don't you subtract 7 from -4 and 23 from 12?

jimthompson5910 (jim_thompson5910):

yes and you divide those differences, which is the order we divide?

jimthompson5910 (jim_thompson5910):

so what do you get when you do the subtractions?

OpenStudy (anonymous):

When you subtract 7 - (-4) = 11 and 23-12= 11 ?

jimthompson5910 (jim_thompson5910):

Good

jimthompson5910 (jim_thompson5910):

Now you divide these differences, what value goes on top?

OpenStudy (anonymous):

When you divide it you get 1

jimthompson5910 (jim_thompson5910):

actually in this case, it doesn't matter, but it's important to know which value goes on top

jimthompson5910 (jim_thompson5910):

yes, the result is 1

jimthompson5910 (jim_thompson5910):

so the slope is 1

jimthompson5910 (jim_thompson5910):

so m = 1 and we can plug that into y = mx+b to get y = 1x+b or just y = x+b

jimthompson5910 (jim_thompson5910):

now we can use any of the two points to find the value of b So say we use (7, 23), this means that x = 7 and y = 23 So y = x+b becomes 23 = 7+b and solving for b gets us: b = 16 So the equation of the line is y = x+16

OpenStudy (anonymous):

I think I have a good understanding, thank you very much

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

But hey, would you solve this problem the same way ? Write an equation in slope-intercept form of the line with an x-intercept of 6 and a y-intercept of 5.

jimthompson5910 (jim_thompson5910):

when they say "x-intercept of 6 and a y-intercept of 5", they mean that this line goes through the points (6, 0) and (0,5)

jimthompson5910 (jim_thompson5910):

so yes, you can use the same technique to solve this one

OpenStudy (anonymous):

So i think I figured it out, but would my answer have 5/6 in it?

jimthompson5910 (jim_thompson5910):

more like -5/6, but you probably lost a minus sign somewhere

jimthompson5910 (jim_thompson5910):

that's the slope of course

OpenStudy (anonymous):

I found my mistake(:

OpenStudy (anonymous):

Thanks for being so helpful.

jimthompson5910 (jim_thompson5910):

yw